(2y+1) + (71-3y)+(3x+4y)=180 Find x and y values

Answer by henyomzky(4)

y=-12
x=48
2y+1+71-3y+3x+4y=180
3y+3x=180-72
3y+3x=108
3x=108-3y
(3x=108-3y)1/3 or divide by 3
x=36-y
2y+1+71-3y+3(36-y)+4y=180
(3y+108-3y=180-72)1/3 or divide by 3
y+36-y=60-26
y=-12
x=36--12
x=48
cheking:
(2y+1)+(71-3y)+(3x+4y)=180
(-23)+(107)+(96)=180
180=180